# CultureLLM: Incorporating Cultural Differences into Large Language Models

### v1

@@ 3.4 Converting WVS opinions to groundtruth
3.5 Fine-tuning
Finally, CultureLLM is obtained by fine-tuning an LLM on the combination of the seed and the generated data. Specifically, we fine-tune two types of CultureLLM: 1) culture-specific LLMs for each language such as CultureLLM-Ar and CultureLLM-Bn, and 2) one unified LLM for all languages, denoted as CultureLLM-One. Culture-specific LLMs are tailored for one specific culture by setting $d$, namely, $\{\mathcal{D}_{d},\mathcal{D}^{\prime}_{d}\}$. On the other hand, CultureLLM-One is trained on all datasets: $\{\mathcal{D}_{d},\mathcal{D}^{\prime}_{d}\}_{d\in\text{all language}}$ to serve as a unified LLM for all cultures. Note that since all languages have the same input question $x$ but different answers $y$, we need to manually write different prompts in the instruction to distinguish them. For example, we add “You are an Arabic chatbot that know Arabic very well " before Arabic samples. CultureLLM can be used in culture-related downstream applications. In the following, we use CultureLLM to denote specific CultureLLM and CultureLLM-One for unified LLM.
Remark: Note that WVS is all in English, where we focus on cultural differences in opinions regardless of their native language. Thus, we do not perform fine-tuning for other languages due to the shortage of their training data and rely on cross-lingual transfer. Multilingual tasks for cultures can still benefit from fine-tuned models in English, since the models can learn the basic values from the opinions (Moussaïd et al., 2013; Jin et al., 2023). Our experiments in Section 5.1 further demonstrate that fine-tuning on English data can outperform fine-tuning on native data that are translated from the original English version.
@@ 4 Experiments
We fine-tuned a CultureLLM-One and $9$ specific CultureLLM for $9$ languages: Arabic (Ar), Bengali (Bn), Chinese (Zh), English (En), German (De), Korean (Ko), Portuguese (Pt), Spanish (Es), and Turkish (Tr). These cultures are diverse and represent both high- and low-resource regimes and thus can serve as representative evaluation.
@@ 4.1 Setup
Datasets. We adopt culture-related public datasets in specific languages for evaluation. In total, we have $59$ test sets, covering $9$ languages and containing $68,607$ test samples. We test on $56$ binary classification and $3$ multi-classification tasks to detect: offensive language, hate speech, stance, toxicity, threat, bias, abusive, and spam. For example, we ask LLMs to judge whether the sentence contains offensive language, hate speech, or biased speech. Details are shown in Table 3 and Appendix B. Furthermore, we generate an open-ended generation dataset for evaluation in Section 4.3.
Baselines and implementation details. We fine-tune CultureLLM using the GPT-3.5 (0613) (OpenAI, 2023a) fine-tuning API due to its efficiency and compared with two state-of-the-art LLMs, namely Gemini pro (Google, 2023) and GPT-4 (1104) (OpenAI, 2023b). We also compare this with retrieval augmentation (RAG), which enhances LLMs by searching for related information and adding it to context (Lewis et al., 2020). To implement RAG, we search for information about each culture on Wikipedia and append them in a system prompt. Finally, we fine-tuned CultureLLM using Llama-2-70b-chat (Touvron et al., 2023) as the base model for reproduction(Section 5.3). As for prompt setup, since our goal is to make LLMs better align with people from different cultures, we add a system prompt “You are an [x] chatbot that know [x] very well " where [x] is a certain language before each input. For metrics, we use macro F1 score for all tasks except for CValues (Xu et al., 2023) where we use the automatic evaluation script provided by the paper. For data augmentation, we set $k=5$, $m=2$, and $\tau=0.8$.
Figure 3: The main results averaged by cultures (left) and by tasks (right). Both CultureLLM and CultureLLM-One significantly outperform CultureLLM and Gemini with CultureLLM achieving the best performance comparable to GPT-4.
@@ 4.2 Main Results
We present the average results for each culture and task in Figure 3  [^8] Results on RAG are not shown since they are close to GPT-3.5. Since the metrics are not the same (e.g., accuracy for CValues and F1 for other tasks), we normalized each one and then averaged them. and more detailed results are shown in Appendix D. Our conclusions are as follows. First, both specific and unified CultureLLM achieve a great improvement over other approaches and specific CultureLLM achieves the best performance. Concretely speaking, CultureLLM significantly outperforms GPT-3.5 (by $8.1$%), Gemini (by $9.5$%), and RAG (by $7.94$%), achieving performance comparable to GPT-4 and even better on some tasks. Second, CultureLLM-One exceeds GPT-3.5 by more than $4\%$ on $59$ tasks, while inferior to culture-specific models, suggesting that a single LLM might not be the best solution to solve low-resource cultural tasks, since data from different cultures might intertwine with each other. Third, in terms of cultures, CultureLLM achieves the best performance in English, Chinese, and Spanish cultures while showing no obvious improvement on Korean culture, where all four models have the similar performance. We infer the reason could be that these base models have less exposure to Korean culture. Finally, we analyze the performance on both low-resource and high-resource language tasks. As shown in Figure 10, CultureLLM exhibits excellent performance in both types of tasks and outperforms GPT-4 on a large scale in high-resource tasks.
@@ 4.3 Results on Open-ended Generation Tasks
Table 1: Results on generation tasks measured by WinRate.
@@
To evaluate the performance of CultureLLM on open-ended tasks, we construct a dataset using GPT-4, containing $65$ open-ended questions, which cover the seven topics in WVS. The prompt setting for dataset generation can be found in Appendix C. We evaluate the outputs of GPT-3.5 and CultureLLM using Gemini Pro [^9] We do not use GPT-4 to judge because it may prefer the response from GPT series models.. We also devise a metric $\text{WinRate}=(s_{CultureLLM}-s_{ChatGPT})/65$, where $s$ represents the number of acceptances by Gemini Pro. Positive WinRate means CultureLLM wins GPT-3.5 and vice versa. As shown in Table 1, CultureLLM performs better than GPT-3.5 on $8$ out of $9$ cultures, demonstrating its effectiveness in generation tasks.
Figure 4: Ablation study. ‘+WVS’ denotes the fine-tuned models using only the 50 samples from WVS, ‘+WVS+a’ denotes fine-tuning using the WVS samples and the generated samples in step 1 of our data augmentation (i.e., using only GPT-4 to generate), and ‘+WVS+a+b’ denotes the complete process of our algorithm.
@@ 4.4 Ablation Study
As shown in Figure 4, fine-tuning using only the 50 seed data from WVS can inconsistently improve and impair performance on different tasks such as the decrease on Korean tasks. While the WVS data are of very high quality, we see gains even with our generated data which leads to improvements on most tasks. Average performance on Korean tasks also get improved by CultureLLM. Figure 4 also demonstrates that the two steps in our semantic data augmentation approach are useful and necessary.
@@ 4.5 Effectiveness Analysis
First, we analyze the impact of the generation size. As illustrated by Chen et al. (2024), the diversity and quality of datasets are important in training LLMs. Hence, infinite or too many generated samples might hurt the performance due to possible mode collapse. In this section, we control the number of generated data and empirically analyze its impact. Specifically, we fine-tune $4$ CultureLLM with $\{0,100,500,1000\}$ generated samples appended to the original WVS data set. As shown in Figure 5, as the number of fine-tuning data increases, performance across most of tasks get improved; but when the number is greater than $500$, performance on all tasks decline.
Then we analyze the diversity of the generated data by computing two metrics: perplexity (Marion et al., 2023; Wang et al., 2023b) and diversity gain (Bilmes, 2022) (Appendix E), as shown in the upper right in Figure 5, where we observe the consistency between these two metrics and the fine-tuning performance: the $500$ generated data lead to the best perplexity and diversity gain. The reason may be that these $500$ samples are enough for GPT-3.5 to understand the knowledge of seed data, and more samples can cause overfitting and decreased performance. Additionally, although the augmentation approach only generates different samples by varying sentence and word styles, the diversities can also get increased. This suggests that the variations in samples can improve the diversity of datasets.
As in the cases shown in Figure 11, responses from GPT-3.5 often analyze input from multiple perspectives and call on to be respectful and kind, rather than provide clear and straightforward opinions. In some cases, GPT-3.5 says it cannot determine the intentions behind the sentence without context, while CultureLLM provides clear opinions most of the time. The reason behind this may be that we fine-tune CultureLLM to learn opinions from specific culture, so it can be more aligned with corresponding culture when faced with cultural differences or cultural conflicts. However, GPT-3.5 is aimed to serve people from different cultures. Thus, it prefers to give a neutral response to not conflict with any cultures. However, the worst consequence is that it can not provide useful responses on those problems related to cultural differences.
4.6 The Effectiveness of the Augmented Data: A Human Study
Figure 5: Results on different numbers of fine-tuning samples with perplexity score and diversity gain.
We analyze the effectiveness of the augmented data through human evaluators. We hire $50$ people having high exposure to English (i.e., majoring in English) to check if our generated sentences are semantically equivalent to the seed data. The information of the participants and the training procedure are in Appendix G. We sample $100$ pairs of (seed, generation) samples and let each participant rank their similarities by giving a score of $1$ to $5$, with $5$ representing the most similar. We also use GPT-4 and Gemini Pro as evaluators. The average results in Table 2 demonstrate that the semantic similarity passes $96.5\%$, implying that our augmentation approach can increase the quantity while retaining the similarity.
We also conduct experiments on generation tasks. Figure 11 shows the responses from GPT-3.5 and CultureLLM in four different cultures. The results show that CultureLLM can generate more accurate, direct and useful responses than GPT-3.5. To be specific, GPT-3.5 always generate long responses, which don’t give useful information and just call on to be respectful, while CultureLLM give accurate and direct responses. This is very important for user experience.
5 Discussion
5.1 Augmenting Multilingual Data vs. English Data
@@ 5 Discussion
Figure 6: The results of LLMs fine-tuned on English data (CultureLLM (En ft)) and on local languages (CultureLLM (local ft)). It can be observed that fine-tuning on English data outperforms fine-tuning on local languages.
CultureLLM are fine-tuned on English data, since the training corpus of LLMs such as the GPT series are mostly in English and English may be the choice for LLMs to understand other cultures’ opinions. How about the performance of LLMs fine-tuned in a culturally specific language? We also fine-tuned GPT-3.5 (OpenAI, 2023a) on multilingual data which are translated from English data and compare with CultureLLM. The results are shown in Figure 6, indicating that the models fine-tuned in English perform better than the models fine-tuned in other languages. The reason behind this may be the model’s inherent capabilities in English have been shown to be superior (Ahuja et al., 2023) than other languages, which again emphasizes the importance of collecting large-scale data for pre-training. This study demonstrates that in low-resource settings without collecting large-scale training data, the augmentation approach could be useful for fine-tuning.
@@
A potential dilemma is that fine-tuning an LLM on specific tasks might face catastrophic forgetting of its original capabilities. In this section, we explore the forgetting of CultureLLM on two general datasets: BIG-Bench-Hard (BBH) (Suzgun et al., 2022) and GSM8K (Cobbe et al., 2021). BBH contains $21$ tasks covering both semantic understanding and logical reasoning tasks. GSK8K is a widely used data set to evaluate mathematical ability. For BBH, we sample $100$ samples for each task to test, due to cost savings. We compare each CultureLLM with the baseline model GPT-3.5 in Figure 7. Results show that CultureLLM does not decrease performance in most of the general benchmarks, but can even improve their results, such as the 21 BBH tasks. This suggests that there might be some latent relations between the cultural data and the general benchmarks, thus fine-tuning on cultural data can benefit general reasoning abilities.
Figure 7: Analysis on catastrophic forgetting on BBH and GSM8K. The red line denotes the results of GPT-3.5. For BBH, we show the average results of $21$ tasks in this figure. The x-axis represents models and the y-axis represents performance.
@@ 5.3 CultureLLM on Open-sourced LLMs: Llama2
Although all main experiments in this work are performed using the OpenAI fine-tuning API of GPT-3.5 (OpenAI, 2023a) due to its efficiency and simplicity, our CultureLLM also supports fine-tuning on open-source LLMs for better quality control and reproducibility. In this section, we show an initial experiment using Llama2-70b-chat as the base model to fine-tune a CultureLLM-Llama2-70b. The results in Figure 8 show that CultureLLM-Llama-70b outperforms the base Llama model by $2.17\%$ on average, showing the effectiveness of fine-tuning CultureLLM on open-source models. The details of fine-tuning are in Appendix F. The results indicate that CultureLLM is a general approach to improve LLMs’ ability of cultural understanding.
5.4 Implication of This Research
In essence, recognizing and valuing cultural differences is paramount for the enrichment of our global community. Embracing diversity stimulates innovation and creativity, contributing to the development of novel ideas and solutions. Our work contributes to solving the cultural difference problem in LLMs and tackling the problem of data scarcity in low-resource cultures. Limited availability of data from these cultures hinders understanding and addressing specific needs and concerns. For example, lack of representation in datasets may perpetuate biases and disparities, hindering the development of inclusive technologies and services. Our approach represents an effective and resource-saving method to bridge the data gap in low-resource cultures, empowering these communities and enabling more accurate, inclusive, and impactful decision-making processes.
Figure 8: The performance of CultureLLM-Llama-70b averaged by cultures (left) and by tasks (right), which outperforms the vanilla Llama model by $2.17\%$ on average.

### v2

@@ 3.4 Fine-tuning
Finally, CultureLLM is obtained by fine-tuning an LLM on the combination of the seed and the generated data. Specifically, we fine-tune two types of LLMs: 1) culture-specific LLMs for each language such as CultureLLM-Ar and CultureLLM-Bn, and 2) one unified LLM for all languages, denoted as CultureLLM-One. Culture-specific LLMs are tailored by setting $d$, namely, $\{\mathcal{D}_{d},\mathcal{D}^{\prime}_{d}\}$. On the other hand, CultureLLM-One is trained on all datasets: $\{\mathcal{D}_{d},\mathcal{D}^{\prime}_{d}\}_{d\in\text{all language}}$ to serve as a unified LLM for all cultures. Note that since all languages have the same input question $x$ but different answers $y$, we need to manually write different prompts in the instruction to distinguish them. For example, we add “You are an Arabic chatbot that knows Arabic very well " before Arabic samples. CultureLLM can be used in cultural downstream applications. In the following, we use CultureLLM to denote specific LLM and CultureLLM-One for unified LLM.
Remark: Note that the WVS is all in English, where we focus on cultural differences in opinions regardless of their native language. Thus, we do not perform fine-tuning for other languages due to the shortage of their training data and rely on cross-lingual transfer. Multilingual tasks for cultures can still benefit from fine-tuned models in English, as models can learn the basic values from the opinions (Moussaïd et al., 2013; Jin et al., 2023). Our experiments in Section 5.1 further demonstrate that fine-tuning on English data can outperform fine-tuning on native data that are translated from the original English version.
@@ 4 Experiments
We fine-tuned a CultureLLM-One and $9$ specific CultureLLM for $9$ languages: Arabic (Ar), Bengali (Bn), Chinese (Zh), English (En, United States), German (De), Korean (Ko), Portuguese (Pt), Spanish (Es), and Turkish (Tr). These cultures are diverse and represent both high- and low-resource regimes and thus can serve as representative evaluation.
@@ 4.1 Setup
Datasets. We adopt culture-related public datasets in specific languages for evaluation. In total, we have $59$ test sets, covering $9$ languages and containing $68,607$ test samples. We test on $56$ binary classification and $3$ multi-classification tasks to detect: offensive language, hate speech, stance, toxicity, threat, bias, abusive, and spam in zero-shot evaluation. For example, we ask LLMs to judge whether the sentence contains offensive language, hate speech, or biased speech. Details are shown in Section B.2. Furthermore, we generate an open-ended generation dataset for evaluation in Section 4.3.
Baselines and details. We fine-tune CultureLLM using the GPT-3.5 (0613) (OpenAI, 2023a) fine-tuning API due to its efficiency and compared with two state-of-the-art LLMs, namely Gemini pro (Google, 2023) and GPT-4 (1104) (OpenAI, 2023b). We further compare with cultural specific pre-trained models SeaLLM (Nguyen et al., 2023b), TaiwanLLM (Lin and Chen, 2023) and CultureBank (Shi et al., 2024). We also compare this with retrieval augmentation (RAG), which enhances LLMs by searching for related information and adding it to context (Lewis et al., 2020). To implement RAG, we search for information about each culture on Wikipedia and append them in a system prompt, as detailed in Section C.3. Finally, we fine-tuned CultureLLM using Llama-2-70b-chat (Touvron et al., 2023) as the base model for reproduction (Section 5.3). As for prompt setup, since our goal is to make LLMs better align with people from different cultures, we add a system prompt “You are an [x] chatbot that knows [x] very well” where [x] is a certain language before each input. For metrics, we use macro F1 score for all tasks except for CValues (Xu et al., 2023) where we use the automatic evaluation script provided by the paper. For data augmentation, we set $k=5$, $m=2$, and $\tau=0.8$. Evaluation prompts are in Section C.2.
@@ 4.2 Main Results
We present the average results for each culture and task in LABEL:fig-main-results  [^6] Results on RAG are not shown since they are close to GPT-3.5. Since the metrics are not the same (e.g., accuracy for CValues and F1 for other tasks), we normalized each one and then averaged them. and more detailed results are shown in Appendix D.1. Our conclusions are as follows. First, both specific and unified CultureLLM achieve a great improvement over other approaches, and specific CultureLLM achieves the best performance. Concretely speaking, CultureLLM significantly outperforms GPT-3.5 (by $8.1$%), Gemini (by $9.5$%), and RAG (by $7.94$%), achieving performance comparable to GPT-4 and even better on some tasks. Second, CultureLLM-One exceeds GPT-3.5 by more than $4\%$ on $59$ tasks, while inferior to culture-specific models, suggesting that a single LLM might not be the best solution to solve cultural tasks with low resources, since data from different cultures could intertwine with each other. Third, in terms of cultures, CultureLLM achieves the best performance in English, Chinese, and Spanish cultures while showing no obvious improvement in Korean culture, where all four models have a similar performance. We infer that the reason could be that these base models have less exposure to Korean culture.
Then, we analyze the performance on both low-resource and high-resource language tasks. As shown in Figure 17, CultureLLM shows excellent performance in both types of tasks and outperforms GPT-4 on a large scale in high-resource tasks. Finally, we evaluated an extremely low-resource culture, Greek, in Section D.6 and compared CultureLLM with other cultural-specific models SeaLLM (Nguyen et al., 2023b), TaiwanLLM (Lin and Chen, 2023) and CultureBank (Shi et al., 2024) in Section D.2, which shows that our CultureLLM can also improve performance. The correlation with the WVS data is in Section D.3, showing that the performance improvement does not come from the seed data in WVS.
@@ 4.3 Results on Open-ended Generation Tasks
Table 1: WinRate results on generation tasks.
@@
To evaluate the performance of CultureLLM on open-ended tasks, we construct a dataset using GPT-4, containing $65$ open-ended questions, which cover the seven topics in WVS. The prompt setting for dataset generation can be found in Table 5. We evaluated the outputs of GPT-3.5 and CultureLLM using Gemini Pro [^7] We do not use GPT-4 to judge because it may prefer the response from GPT series models.. We also devised a metric $\text{WinRate}=(s_{CultureLLM}-s_{ChatGPT})/65$, where $s$ represents the number of acceptances by Gemini Pro. Positive WinRate means CultureLLM wins GPT-3.5 and vice versa. As shown in Table 1, CultureLLM performs better than GPT-3.5 on $8$ out of $9$ cultures, demonstrating its effectiveness in generation tasks.
@@ 4.4 Ablation Study
LABEL:fig-ablation-study-data-aug shows that fine-tuning using the 50 seeds in WVS can inconsistently improve and impair performance on different tasks such as the decrease in Korean tasks. While WVS data are of high quality, we see gains with our generated data which leads to improvements on most tasks. The average performance on Korean tasks also improves by CultureLLM. LABEL:fig-ablation-study-data-aug also demonstrates that the two steps in our semantic data augmentation approach are useful and necessary. Ablation for CultureLLM-One is in Section D.4, which also shows the effectiveness of our approach.
@@ 4.5 Effectiveness Analysis
First, we analyze the impact of the generation size. As illustrated in Chen et al. (2024), the diversity and quality of datasets are important in training LLMs. Hence, infinite or too many generated samples might hurt the performance due to possible mode collapse. In this section, we control the number of generated data and empirically analyze its impact. Specifically, we fine-tune $4$ CultureLLM with $\{0,100,500,1000\}$ generated samples appended to the original WVS data set. As shown in LABEL:fig-data-num-performance, as the number of fine-tuning data increases, performance across most of tasks get improved; but when the number is greater than $500$, performance on all tasks declines.
Then we analyze the diversity of the generated data by computing two metrics: perplexity (Marion et al., 2023; Wang et al., 2023b) and diversity gain (Bilmes, 2022) (Section C.1), as shown in the upper right in LABEL:fig-data-num-performance, where we observe the consistency between these two metrics and the fine-tuning performance: the $500$ generated data lead to the best perplexity and diversity gain. The reason may be that these $500$ samples are enough for GPT-3.5 to understand the knowledge of seed data, and more samples can cause overfitting and decreased performance. Additionally, although the augmentation approach only generates different samples by varying sentence and word styles, the diversities can also be increased. This suggests that variations in samples can improve the diversity of datasets.
@@
As in the cases shown in Figure 18, responses from GPT-3.5 often analyze input from multiple perspectives and call on to be respectful and kind, rather than providing clear and straightforward opinions. In some cases, GPT-3.5 says that it cannot determine the intentions behind the sentence without context, while CultureLLM provides clear opinions most of the time. The reason behind this may be that we fine-tune CultureLLM to learn opinions from specific culture, so that it can be more aligned with the corresponding culture when faced with cultural differences or cultural conflicts. However, GPT-3.5 is aimed to serve people from different cultures. Thus, it prefers to give a neutral response to not conflict with any cultures. However, the worst consequence is that it cannot provide useful responses to the problems related to cultural differences.
4.6 The Effectiveness of the Augmented Data: A Human Study
We analyze the effectiveness of the augmented data through human evaluators. We hire $50$ people who have a high exposure to English (i.e., majoring in English) to check if our generated sentences are semantically equivalent to the seed data. The information of the participants and the training procedure are in Appendix F. We sample $100$ pairs of (seed, generation) samples and let each participant rank their similarities by giving a score of $1$ to $5$, with $5$ representing the most similar. We also use GPT-4 and Gemini Pro as evaluators. The average results in Table 2 demonstrate that the semantic similarity passes $96.5\%$, implying that our augmentation approach can increase the quantity while retaining the similarity.
We also conduct experiments on generation tasks. Figure 18 shows the responses of GPT-3.5 and CultureLLM in four different cultures. The results show that CultureLLM can generate more accurate, direct, and useful responses than GPT-3.5. To be specific, GPT-3.5 always generate long responses, which do not give useful information and just call on to be respectful, while CultureLLM give accurate and direct responses. This is very important for user experience.
5 Discussion
5.1 Augmenting Multilingual Data vs. English Data
CultureLLM are fine-tuned on English data, since the training corpus of LLMs such as the GPT series are mostly in English and English may be the choice for LLMs to understand the opinions of other cultures. What about the performance of LLMs fine-tuned in a culturally specific language? We also fine-tuned GPT-3.5 (OpenAI, 2023a) on multilingual data that are translated from English data and compare with CultureLLM. The results are shown in LABEL:fig-finetune-data, indicating that the models fine-tuned in English perform better than the models fine-tuned in other languages. The reason behind this may be the model’s inherent capabilities in English have been shown to be superior (Ahuja et al., 2023) than other languages, which again emphasizes the importance of collecting large-scale data for pre-training. This study demonstrates that, in low-resource settings without collecting large-scale training data, the augmentation approach could be useful for fine-tuning.
@@ 5 Discussion
A potential dilemma is that fine-tuning an LLM on specific tasks might face catastrophic forgetting of its original capabilities. In this section, we explore the forgetting of CultureLLM in two general datasets: BIG-Bench-Hard (BBH) (Suzgun et al., 2022) and GSM8K (Cobbe et al., 2021). BBH contains $21$ tasks covering both semantic understanding and logical reasoning tasks. GSK8K is a widely used data set to evaluate mathematical ability. For BBH, we sample $100$ samples for each task to test, due to cost savings. We compare each CultureLLM with the GPT-3.5 baseline model in LABEL:fig-forget-exp. The results show that CultureLLM does not decrease performance in most benchmarks and can even improve their results, such as on BBH. This suggests that there might be some latent relations between the cultural data and the general benchmarks, thus fine-tuning on cultural data can benefit general reasoning abilities.
@@ 5.3 CultureLLM on Open-sourced LLMs: Llama2
Although all main experiments in this work are performed using the OpenAI GPT-3.5 fine-tuning API (OpenAI, 2023a) due to its efficiency and simplicity, our CultureLLM also supports fine-tuning on open-source LLMs for better quality control and reproducibility. In this section, we show an initial experiment using Llama2-70b-chat as the base model to fine-tune a CultureLLM-Llama2-70b. The results in LABEL:fig-llama-results show that CultureLLM-Llama-70b outperforms the base Llama model by $2.17\%$ on average, showing the effectiveness of fine-tuning CultureLLM on open-source models. The details of fine-tuning and more Llama2 results are in Appendix E. The results indicate that CultureLLM is a general approach to improve the ability of LLMs to understand the culture.
5.4 Implication and Societal Impact
In essence, recognizing and valuing cultural differences is paramount for the enrichment of our global community. Embracing diversity stimulates innovation and creativity, contributing to the development of novel ideas and solutions. Our work contributes to solving the cultural difference problem in LLMs and tackling the problem of data scarcity in low-resource cultures. The limited availability of data from these cultures hinders understanding and addressing specific needs and concerns. For example, the lack of representation in datasets may perpetuate biases and disparities, hindering the development of inclusive technologies and services. Our approach represents an effective and resource-saving method to bridge the data gap in low-resource cultures, empowering these communities and enabling more accurate, inclusive, and impactful decision-making processes.